# FeNNol: an Efficient and Flexible Library for Building Force-field-enhanced Neural Network Potentials

### v1

@@ I Introduction
However, the advent of these new models and architectures has brought its own set of challenges. Prominently, numerical implementations of these architectures are currently scattered in many packages provided by the authors of the methods . This makes a direct comparison between the models a difficult task, in particular with respect to computational efficiency. A few existing libraries provide a wider scope and more modular implementations such as SchnetPack Schütt et al. (2023), MLAtom Dral et al. (2024) or deepmd-kit Zeng et al. (2023) but they are still limited in the range of models offered to the user. Furthermore, designing models for non-standard workflows usually requires custom implementations, thus imposing barriers at entry for non-expert users. Examples of such applications include, for example, fitting distributed multipoles or a decomposition of the energy, or feeding force-field terms with ML-generated parameters, as was done in our previously proposed FeNNix model Plé, Lagardère, and Piquemal (2023).
@@
To address these challenges, we present FeNNol, a new Python library designed for building, training, and running machine-learning potentials, with a particular focus on physics-enhanced neural networks. FeNNol provides a flexible and modular system that allows users to easily build custom models, allowing for example the combination of state-of-the-art atomic embeddings with ML-parameterized physical interaction terms, without the need for explicit programming. FeNNol leverages the Jax framework Bradbury et al. (2018); DeepMind et al. (2020) and its just-in-time compilation capabilities, and provides a collection of efficient and configurable modules that can be composed to form complex models.
@@
The paper is structured as follows: in section II, we delve into the software architecture of FeNNol, detailing its design principles and the underlying technologies it leverages. In section III, we introduce a novel multi-paradigm atomic embedding which serves as a demonstration of the highly configurable nature of FeNNol modules. In section IV, we describe FeNNol’s training system that enables users to define complex models and train them on generic tasks. In section V, we present the different ways to use FeNNol models in molecular dynamics simulation. We also demonstrate the performance of FeNNol’s models and native MD engine by showing that our implementation of the popular ANI-2x model Devereux et al. (2020) reaches simulation speeds close to the optimized GPU-accelerated (GPU=Graphics Processing Unit) Tinker-HP implementation of the AMOEBA force-field Adjoua et al. (2021). Finally, in section VI, we provide concrete examples of model training with FeNNol, demonstrating how one can easily build non-standard models and how state-of-the-art results can be achieved within a few hours of training on a commodity GPU.
@@ II Software Architecture
A global description of the software architecture and data flows is provided in figure 1. In the following subsections, we describe the different modules available to build models, starting with the preprocessing modules that handle operations on neighbour lists and following with the main modules that handle the bulk of the computation and are divided in a few categories: atomistic embeddings, chemical and radial encodings, physics modules, neural networks and operation modules. We will then describe how one can extend FeNNol by registering custom modules. Finally, we provide an example of how to build a full model with the FENNIX constructor which is the main entry point for model building.
@@
Figure 1: Architecture and data flow of the FeNNol library. Black blocks represent functional modules, green arrows represent model operations and lue arrows represent data flows.
II.1 Preprocessing
@@
As explained earlier, a preprocessing pipeline that is not compiled is necessary to build neighbour lists without assuming a maximum size in advance. Preprocessing modules thus handle operations on neighbour lists – or more generally molecular graphs – such as building, filtering, padding or extending.
@@
The main preprocessing module is the GraphGenerator which constructs a molecular graph given a cutoff distance, a set of atom coordinates, and optionally cell vectors for periodic boundary conditions. To avoid padding the array of coordinates when a batch of different molecular structures is provided, we chose to handle a concatenated array of all the atoms, accompanied by an array keeping track of the batch index of each atom. A molecular graph consists a minima of two arrays of atom indices: the source and destination of each edge in the molecular graph. An edge is assigned between two atoms when they have the same batch index and the distance between them is lower than the cutoff distance (the source and destination arrays thus form the neighbour list). Note that in the current implementations, all graphs are directed (meaning that each pair of atoms appears twice in the neighbour list, with source and destination atoms reversed). If cell vectors are provided edges are assigned for each pair of images within the cutoff distance (with at least one of the atom in the main periodic cell) or according to the minimum image convention if required by the user. When periodic boundary conditions (PBCs) are applied, an array containing the shift (in units of cell vectors) to apply to the vector connecting the source and destination atoms is added to the graph. In the current implementation, neighbour lists are obtained by first computing all distances between pairs of atoms (that have the same batch index) and then filtering to the cutoff distance. This implementation is efficient for small number of atoms and with GPU acceleration (and using a neighbour list skin in MD simulations to compute all distances less often), we found that systems up 20,000 atoms do not suffer too much from the quadratic scaling. Future releases will use cell lists to improve the scaling of neighbour lists construction. Finally, when the model uses long-range interactions that require Ewald summationsEwald (1921), the user can configure the GraphGenerator module to construct vectors from the reciprocal lattice and filter them according to a magnitude cutoff.
@@ II.2 Main Modules
Atomic embedding modules process the molecular graph to compute descriptors of the atoms environments that can later be used by other modules (such as neural networks) to compute properties of the system (for example the energy). Embeddings are usually of atom-wise or pair-wise nature and can be either local (that only process directly connected atoms), message-passing (that capture longer-range interactions via iterative exchanges of information with neighbouring atoms) or completely non-local (for example global embeddings such as Coulomb matrices Rupp et al. (2012) or non-local atom-wise like in the LODE framework Grisafi and Ceriotti (2019); Grisafi, Nigam, and Ceriotti (2021) or the attention mechanism in SpookyNet Unke et al. (2021)). FeNNol provides various embeddings from the literature: local models such as ANI AEVs Smith, Isayev, and Roitberg (2017), Allegro Musaelian et al. (2023), EEACSF Eckhoff and Reiher (2023) and DeepPot-SE Zhang et al. (2018) and message-passing models such as AIMNet Zubatyuk et al. (2019), HIP-NN Lubbers, Smith, and Barros (2018), PaiNN Schütt, Unke, and Gastegger (2021) and NewtonNet Haghighatlari et al. (2022). When necessary, E(3)-equivariant operations (such as tensor products) are performed using the e3nn-jax Geiger and Smidt (2022) library. In some simple cases, we implemented equivariant operations from scratch to increase performance (for example, the filtered tensor product proposed in ref. Kozinsky et al., 2023). We also implemented a new multi-paradigm embedding denoted as CRATE that will be described in section III.
@@ II.3 Extensions with custom modules
When importing the FeNNol library, all the flax modules in the source code with the FID (Fennol IDentifier) attribute are automatically registered as FeNNol modules and available when building a model. The user can also manually register a module without modifying the source code by providing their custom module and an FID to the register_fennix_module function before building or loading a model.
@@ II.4 Building a FeNNix model
In FeNNol, a model is constructed by calling the FENNIX constructor which most important arguments are two ordered dictionaries: one that lists the preprocessing modules and their properties and another one for the main modules with their properties:
@@
The model automatically builds a primary graph with a cutoff radius specified by the cutoff argument. This primary graph can optionally customized using the preprocessing parameters. The argument rng_key is a key for the random number generator that is used to initialize the model parameters. The energy_terms argument defines which keys in the final system state will be used as energy components and summed to form the total energy. If only one graph is required (i.e. no filtering and no angle information needed), the preprocessing argument can be omitted. The modules dictionary has the following structure:
@@
A flow diagram of the execution of a ANI-like model built with FeNNol is provided in figure 2. The input is built from chemical species (atomic numbers) and coordinates. In case where multiple structures are provided (batching), all the atoms are concatenated in the species and coordinates arrays and an additional natoms array is provided containing the number of atoms of each structure in the batch and is used to compute the batch index of each atom. The first modules on the left are the preprocessing modules that build the neighbour list (with a cutoff of 5 Å), filters it to a new graph with a cutoff of 3.5 Å  and finally extend this graph with indices of pairs of neighbours forming angles that will be used to construct the environment vector in the ANI_AEV module. The blue arrows indicate graph processing modules that are automatically added by the corresponding preprocessing modules. As explained earlier, these processing modules compute distances and angles using provided edge indices that will be automatically differentiated (green box in the diagram) to obtain atomic forces. The next two modules are the user-defined main modules (here ANI_AEV that computes the atomic environment vectors of the ANI model Smith, Isayev, and Roitberg (2017) and ChemicalNet that applies a species-specialized neural network to the AEVs to obtain atomic energies). The last automatically differentiated block simply sums the energy contributions of each atom to the corresponding structure’s total energy. The final module adds atomic forces obtained from the jax.grad transform to the output.
@@ III The "CRATE" multi-paradigm embedding
By selecting specific resources and interaction layers, the CRATE embedding can be made to operate similarly as well-established architectures. For example, selecting radial and angular information with only one interaction layer, the model behaves like the ANI or BPNN architectures. Keeping the same resources and adding multiple message-passing layers leads to an architecture that resembles AIMNet Zubatyuk et al. (2019). Finally, if we select E(3)-equivariant resources, refine them with local atom-wise tensor products and combine them using message-passing interactions, we obtain a model close to the recently proposed MACE Batatia et al. (2022). Each interaction layer can be configured separately so that many combinations are possible and one can tailor the architecture for their data and computational efficiency requirements.
@@ IV Model Training
FeNNol provides a basic training system that should handle most simple model training tasks. The input to the training program is a structured file (either yaml, json or toml) that describes both the training parameters and the model structure. The latter is the same as presented in section II.4. The training parameters can be subdivided in three categories: general parameters, optimizer and schedule parameters and loss function definition.
@@
Multiple loss types are provided such as mean absolute error, log hyperbolic cosine, negative log-likelihood or CRPS for ensemble training Kellner and Ceriotti (2024) or evidential losses Amini et al. (2020); Meinert and Lavin (2021). The loss type can be set for all components at once using the default_loss_type parameter. The per_atom directive only applies to system-wise properties and divides the loss by the number of atoms. This can be useful when training on systems of widely varying sizes.
@@
When training a model on a multi-component loss function, it may be useful to define a multi-stage training. As a concrete example, we can mention the training of the FENNIX-OP1 model Plé, Lagardère, and Piquemal (2023) which required to first train on short-range energies and partial charges, and then training again with long-range energy contributions activated. Another application is transfer learning where we would for example pre-train a model on a large DFT dataset and than transfer the model to CCSD(T) accuracy on a smaller dataset. In order to facilitate the reproducibility of such multi-stage, it can be fully defined in a single input file to the FeNNol training system. The following listing shows an example of a possible transfer-learning task:
@@
In its current version, FeNNol only accepts a simple but specific structure for datasets. The root of the dataset should be a dictionary with at least the two entries training and validation that are both lists containing the molecular configurations (one per element in the list) that will be used to train and validate the model respectively. Each molecular configuration is itself a dictionary which should contain the entries species (an array of atomic numbers of the atoms in the configuration) and coordinates (the array of 3D coordinates of the atoms), as well as all the keys used as reference in the definition of the training loss. The dataset must be formatted by the user and can be provided in multiple formats. The simplest is to use the generic pickle format to save the dataset from a python script. This format will be completely loaded in memory before training and loading batches is thus very fast. This format is best suited for small to moderately-large datasets as it must fit in memory and initial loading times can be very long for large datasets. In most applications, it should however be the simplest format and we found it practical even for quite large datasets (for example the ANI-2x dataset Devereux et al. (2020) with almost ten million entries). For larger datasets, one can provide a hdf5 file or a sqlite database that can be queried without loading it in memory completely. These format may however have slower access times.
@@ V Running Molecular Dynamics simulations
FeNNol offers multiple ways of running molecular dynamics (MD) simulations with FeNNix models. One can either use custom python scripts, the provided calculator for the Atomic Simulation Environment Larsen et al. (2017) (ASE) package, the Tinker-HP MD engine  Lagardère et al. (2018); Adjoua et al. (2021) via the Deep-HP interface Inizan et al. (2022) or FeNNol’s native MD engine.
@@ V.2 ASE calculator
FeNNol provides a FENNIXCalculator class that can be used with the Atomic Simulation Environment Larsen et al. (2017) (ASE) package. This enables the use of all the advanced functionalities provided by ASE such as geometry optimization, phonon calculations and various MD engines in a very user-friendly manner. For example, a geometry optimization followed by a short MD simulation with a FeNNix model can be obtained with the following simple code:
@@ V.3 Deep-HP interface
We modified the Deep-HP interface introduced in ref. Inizan et al., 2022 to enable the use of FeNNix models as energy components in Tinker-HP. This interface has multiple benefits. First, it allows the use of the different integrators, thermostats, barostats and accelerated sampling methods provided by Tinker-HP Lagardère et al. (2018); Adjoua et al. (2021). Efficient methods for simulating NQEs are also readily available with the Quantum-HP module Plé et al. (2023). Second, it enables the parallelization of FeNNix models over multiple GPUs and multiple nodes, opening up simulations of very large systems with ML models. Note that currently, multi-GPU parallelization is only safe with purely local models such as ANI or Allegro. Finally, FeNNix models can be combined with traditional force fields, either via an ML/MM embedding scheme like in refs. Inizan et al., 2022; Lahey and Rowley, 2020; Galvelis et al., 2023, or as energy components in a hybrid model as in refs. Illarionov et al., 2023; Wang et al., 2024b.
As in ref. Inizan et al., 2022, we directly convert GPU pointers allocated from Tinker-HP (for example coordinates or neighbour lists) to Jax arrays via the DLPack protocol in order to minimize data movements and CPU/GPU communications. Note that the Tinker-HP/FeNNol interface is currently in early stages of development and will be refined and optimized in future works.

### v2

@@ I Introduction
However, the advent of these new models and architectures has brought its own set of challenges. Prominently, numerical implementations of these architectures are currently scattered in many packages provided by the authors of the methods . This makes a direct comparison between the models a difficult task, in particular with respect to computational efficiency. A few existing libraries provide a wider scope and more modular implementations such as SchnetPack Schütt et al. (2023), MLAtom Dral et al. (2024) or deepmd-kit Zeng et al. (2023) but they are still limited in the range of models offered to the user. Furthermore, designing models for non-standard workflows usually requires custom implementations, thus imposing barriers at entry for non-expert users. Examples of such applications include, for example, fitting distributed multipoles Bereau, Andrienko, and Von Lilienfeld (2015); Glick et al. (2021); Thurlemann, Boselt, and Riniker (2022) or a decomposition of the energy, or feeding force-field terms with ML-generated parameters, as was done in our previously proposed FeNNix model Plé, Lagardère, and Piquemal (2023).
@@
To address these challenges, we present FeNNol, a new Python library designed for building, training, and running atomistic machine-learning models, with a particular focus on physics-enhanced neural network potentials. FeNNol provides a flexible and modular system that allows users to easily build custom models, allowing for example the combination of state-of-the-art atomic embeddings with ML-parameterized physical interaction terms, without the need for explicit programming. FeNNol leverages the Jax framework Bradbury et al. (2018); DeepMind et al. (2020) and its just-in-time compilation capabilities, and provides a collection of efficient and configurable modules that can be composed to form complex models.
@@
The paper is structured as follows: in section II, we describe FeNNol’s software architecture, detailing its design principles and the underlying technologies it leverages. We focus in particular on FeNNol "modules" which are at the core of model building. In section III, we introduce a novel multi-paradigm atomic embedding which serves as a demonstration of the highly configurable nature of FeNNol modules. In section IV, we describe FeNNol’s training system that enables users to define complex models and train them on generic tasks. In section V, we present the different ways to use FeNNol models in molecular dynamics simulation. We then demonstrate the performance of FeNNol’s models and native MD engine by showing that our implementation of the popular ANI-2x model Devereux et al. (2020) reaches simulation speeds close to the optimized GPU-accelerated (GPU=Graphics Processing Unit) Tinker-HP implementation of the AMOEBA force-field Adjoua et al. (2021). We also showcase our interface with Deep-HP Jaffrelot Inizan et al. (2023) that enables efficient NNP simulations of systems reaching half a million atoms. Finally, in section VI, we provide concrete examples of model training with FeNNol, demonstrating how one can easily build non-standard models and how state-of-the-art results can be achieved within a few hours of training on a commodity GPU.
@@ II Software Architecture
A global description of the software architecture and data flows is provided in figure 1. In the following subsections, we describe the different modules available to build models, starting with the preprocessing modules that handle operations on neighbour lists and following with the computation modules that handle the bulk of models’ logic and are divided in a few categories: atomistic embeddings, chemical and radial encodings, physics modules, neural networks and operation modules. We will then describe how one can extend FeNNol by registering custom modules. Finally, we provide an example of how to build a full model with the FENNIX constructor which is the main entry point for model building.
@@
Figure 1: Architecture and data flow of the FeNNol library. Black blocks represent functional modules, green arrows represent model operations and blue arrows represent data flows.
II.1 Preprocessing Modules
@@
As explained earlier, a preprocessing pipeline that is not compiled is necessary to build neighbour lists without assuming a maximum size in advance. The preprocessing modules thus handle operations on neighbour lists – or more generally molecular graphs – such as building, filtering, padding or extending.
@@
The main preprocessing module is the GraphGenerator which constructs a molecular graph given a cutoff distance, a set of atom coordinates, and optionally cell vectors for periodic boundary conditions. To avoid padding the array of coordinates when a batch of different molecular structures is provided, we chose to handle a concatenated array of all the atoms, accompanied by an array keeping track of the batch index of each atom. A molecular graph consists a minima of two arrays of atom indices: the source and destination of each edge in the molecular graph. An edge is assigned between two atoms when they have the same batch index and the distance between them is lower than the cutoff distance. Note that in the current implementations, all graphs are directed (meaning that each pair of atoms appears twice in the neighbour list, with source and destination atoms reversed). If cell vectors are provided, edges are assigned for each pair of images within the cutoff distance (with at least one of the atom in the main periodic cell) or according to the minimum image convention if required by the user. When periodic boundary conditions (PBCs) are applied, an array containing the shift (in units of cell vectors) to apply to the vector connecting the source and destination atoms is added to the graph. In the current implementation, neighbour lists are obtained by first computing all distances between pairs of atoms (that have the same batch index) and then filtering to the cutoff distance. This implementation is efficient for small number of atoms and, with GPU acceleration (and using a neighbour list skin in MD simulations to compute all distances less often), we found that systems up 20,000 atoms do not suffer too much from the quadratic scaling. Future releases will use cell lists to improve the scaling of neighbour lists construction. Finally, when the model uses long-range interactions that require Ewald summationsEwald (1921), the user can configure the GraphGenerator module to construct vectors from the reciprocal lattice and filter them according to a magnitude cutoff.
@@ II.2 Computation Modules
Atomic embedding modules process the molecular graph to compute descriptors of the atoms’ environments that can later be used by other modules (such as neural networks) to compute properties of the system (for example the energy). Embeddings are usually of atom-wise or pair-wise nature and can be either local (that only process directly connected atoms), message-passing (that capture longer-range interactions via iterative exchanges of information with neighbouring atoms) or completely non-local (for example global embeddings such as Coulomb matrices Rupp et al. (2012) or non-local atom-wise like in the LODE framework Grisafi and Ceriotti (2019); Grisafi, Nigam, and Ceriotti (2021) or the attention mechanism in SpookyNet Unke et al. (2021)). FeNNol provides various embeddings from the literature: local models such as ANI AEVs Smith, Isayev, and Roitberg (2017), Allegro Musaelian et al. (2023), EEACSF Eckhoff and Reiher (2023) and DeepPot-SE Zhang et al. (2018) and message-passing models such as AIMNet Zubatyuk et al. (2019), HIP-NN Lubbers, Smith, and Barros (2018), PaiNN Schütt, Unke, and Gastegger (2021) and NewtonNet Haghighatlari et al. (2022). When necessary, E(3)-equivariant operations (such as tensor products) are performed using the e3nn-jax Geiger and Smidt (2022) library. In some simple cases, we implemented equivariant operations from scratch to increase performance (for example, the filtered tensor product proposed in ref. Kozinsky et al., 2023). We also implemented a new multi-paradigm embedding denoted as CRATE that will be described in section III.
@@ II.3 Extensions with custom modules
When importing the FeNNol library, all the flax modules in the source code with the FID (Fennol IDentifier) attribute are automatically registered as FeNNol modules and available when building a model. FeNNol also automatically registers modules from the Python files found in repositories listed in the environment variable FENNOL_MODULES_PATH. The user can also manually register a module after initialization by providing their custom module and an FID to the register_fennix_module function before building or loading a model.
@@ II.4 Building a FeNNix model
In FeNNol, a model is constructed by calling the FENNIX constructor which most important arguments are two ordered dictionaries: one that lists the preprocessing modules and their properties and another one for the computation modules with their properties:
@@
The model automatically builds a primary graph with a cutoff radius specified by the cutoff argument. This primary graph can optionally be customized using the preprocessing parameters. The argument rng_key is a key for the random number generator that is used to initialize the model parameters. The energy_terms argument defines which keys in the final system state will be used as energy components and summed to form the total energy. If only one graph is required (i.e. no filtering and no angle information needed), the preprocessing argument can be omitted. The modules dictionary has the following structure:
@@
A flow diagram of the execution of a ANI-like model built with FeNNol is provided in figure 2. The input is built from chemical species (atomic numbers) and coordinates. In case where multiple structures are provided (batching), all the atoms are concatenated in the species and coordinates arrays and an additional natoms array is provided containing the number of atoms of each structure in the batch and is used to compute the batch index of each atom. The first modules on the left are the preprocessing modules that build the neighbour list (with a cutoff of 5 Å), filter it to a new graph with a cutoff of 3.5 Å  and finally extend this graph with indices of pairs of neighbours forming angles that will be used to construct the environment vector in the ANI_AEV module. The blue arrows indicate graph processing modules that are automatically added by the corresponding preprocessing modules. As explained earlier, these processing modules compute distances and angles using provided edge indices that will be automatically differentiated (green box in the diagram) to obtain atomic forces. The next two modules are the user-defined main modules (here ANI_AEV that computes the atomic environment vectors of the ANI model Smith, Isayev, and Roitberg (2017) and ChemicalNet that applies a species-specialized neural network to the AEVs to obtain atomic energies). The last automatically differentiated block simply sums the energy contributions of each atom to the corresponding structure’s total energy. The final module adds atomic forces obtained from the jax.grad transform to the output.
@@ III The "CRATE" multi-paradigm embedding
By selecting specific resources and interaction layers, the CRATE embedding can be made to operate similarly to well-established architectures. For example, selecting radial and angular information with only one interaction layer, the model behaves like the ANI or BPNN architectures. Keeping the same resources and adding multiple message-passing layers leads to an architecture that resembles AIMNet Zubatyuk et al. (2019). Finally, if we select E(3)-equivariant resources, refine them with local atom-wise tensor products and combine them using message-passing interactions, we obtain a model close to the recently proposed MACE Batatia et al. (2022). Each interaction layer can be configured separately so that many combinations are possible and one can tailor the architecture for their data and computational efficiency requirements.
@@ IV Model Training
FeNNol provides a basic training system that should handle most simple model training tasks. This training system can be invoked with the command fennol_train that is installed with the FeNNol Python package. The input to the training program is a structured file (either yaml, json or toml) that describes both the training parameters and the model structure. The latter is the same as presented in section II.4. The training parameters can be subdivided in three categories: general parameters, optimizer and schedule parameters and loss function definition.
@@
Multiple loss types are provided such as means squared error, mean absolute error, log hyperbolic cosine, negative log-likelihood or CRPS for ensemble training Kellner and Ceriotti (2024) or evidential losses Amini et al. (2020); Meinert and Lavin (2021). The loss type can be set for all components at once using the default_loss_type parameter. The per_atom directive only applies to system-wise properties and divides the loss by the number of atoms. This can be useful when training on systems of widely varying sizes.
@@
When training a model on a multi-component loss function, it may be useful to define a multi-stage training. As a concrete example, we can mention the training of the FENNIX-OP1 model Plé, Lagardère, and Piquemal (2023) which required to first train on short-range energies and partial charges, and then training again with long-range energy contributions activated. Another application is transfer learning where we would for example pre-train a model on a large DFT dataset and than transfer the model to CCSD(T) accuracy on a smaller dataset. In order to facilitate the reproducibility of such multi-stage training, it can be fully defined in a single input file to the FeNNol training system. The following listing shows an example of a possible transfer-learning task:
@@
In its current version, FeNNol only accepts a simple but specific structure for datasets. The root of the dataset should be a dictionary with at least the two entries training and validation that are both lists containing the molecular configurations (one per element in the list) that will be used to train and validate the model respectively. Each molecular configuration is itself a dictionary which should contain the entries species (an array of atomic numbers of the atoms in the configuration) and coordinates (the array of 3D coordinates of the atoms), as well as all the keys used as reference in the definition of the training loss. The dataset must be formatted by the user and can be provided in multiple formats. The simplest is to use the generic pickle format to save the dataset from a python script. This format will be completely loaded in memory before training and loading batches is thus very fast. This format is best suited for small to moderately-large datasets as it must fit in memory and initial loading times can be very long for large datasets. In most applications, it should however be the simplest format and we found it practical even for quite large datasets (for example the ANI-2x dataset Devereux et al. (2020) with almost ten million entries). For larger datasets, one can provide a hdf5 file or a sqlite database that can be queried without loading it in memory completely. These formats may however have slower access times.
@@ V Running Molecular Dynamics simulations
FeNNol offers multiple ways of running molecular dynamics (MD) simulations with FeNNix models. One can either use custom python scripts, the provided calculator for the Atomic Simulation Environment Larsen et al. (2017) (ASE) package, the Tinker-HP MD engine  Lagardère et al. (2018); Adjoua et al. (2021) via the Deep-HP interface Jaffrelot Inizan et al. (2023) or FeNNol’s native MD engine.
@@ V.2 ASE calculator
FeNNol provides a FENNIXCalculator class that can be used with the Atomic Simulation Environment Larsen et al. (2017) (ASE) package. This enables the use of all the advanced functionalities provided by ASE such as geometry optimization, phonon calculations and various MD algorithms in a very user-friendly manner. For example, a geometry optimization followed by a short MD simulation with a FeNNix model can be obtained with the following simple code:
@@ V.3 Deep-HP interface
We modified the Deep-HP interface introduced in ref. Jaffrelot Inizan et al., 2023 to enable the use of FeNNix models as energy components in Tinker-HP. This interface has multiple benefits. First, it allows the use of the different integrators, thermostats, barostats and accelerated sampling methods provided by Tinker-HP Lagardère et al. (2018); Adjoua et al. (2021). Efficient methods for simulating NQEs are also readily available with the Quantum-HP module Plé et al. (2023). Second, it enables the parallelization of FeNNix models over multiple GPUs and multiple nodes, opening up simulations of very large systems with ML models. Note that currently, multi-GPU parallelization is only safe with purely local models such as ANI or Allegro. Finally, FeNNix models can be combined with traditional force fields, either via an ML/MM embedding Jaffrelot Inizan et al. (2023); Lahey and Rowley (2020); Galvelis et al. (2023), or as energy components in a hybrid model Illarionov et al. (2023); Wang et al. (2024b).
As in ref. Jaffrelot Inizan et al., 2023, we directly convert GPU pointers allocated from Tinker-HP (for example coordinates or neighbour lists) to Jax arrays via the DLPack protocol in order to minimize data movements and CPU/GPU communications. Note that the Tinker-HP/FeNNol interface is currently in early stages of development and will be refined and optimized in future works.